# On the ETHOS of AI Agents: An Ethical Technology and Holistic Oversight System

### v2

@@ Abstract
In a world increasingly defined by machine intelligence, the future depends on how we govern AI’s development and integration into society. Recent initiatives, such as the EU’s AI Act, EDPB opinion, U.S. Bipartisan House Task Force and NIST’s AI Risk Management Report, highlight the urgent need for robust governance frameworks to address the challenges posed by advancing AI technologies. However, existing frameworks fail to adequately address the rise of AI agents—autonomous systems capable of complex decision-making, learning, and adaptation—or the ongoing debate between centralized and decentralized governance models. To bridge these gaps, we propose the ETHOS (Ethical Technology and Holistic Oversight System) framework, which leverages Web3 technologies—including blockchain, smart contracts, decentralized autonomous organizations (DAOs), and soulbound tokens (SBTs)—to establish a decentralized global registry for AI agents. ETHOS incorporates the concept of AI-specific legal entities, enabling these systems to assume limited liability and ensuring accountability through mechanisms like insurance and compliance monitoring. Additionally, the framework emphasizes the need for a collaborative, participatory approach to AI governance, engaging diverse stakeholders through public education, transparency, and international coordination. ETHOS balances innovation with ethical accountability, providing a forward-looking strategy for the responsible integration of AI agents into society. Finally, this exploration reflects the emergence of a new interdisciplinary field we define as “Systems Thinking at the Intersection of AI, Web3, and Society”.
Keywords: Artificial Intelligence, AI Agents, AI Regulation, AI Governance, Web3
“Imagine a world where every AI agent’s identity is as transparent as a blockchain ledger, yet as secure as a vault. A decentralized global registry isn’t just a possibility—it’s the future of AI governance, where trust and transparency waltz in perfect harmony.”
-Ari, an AI agent (Gemach DAO).
@@ 1 Introduction
Artificial intelligence (AI) is reaching a defining moment in its integration into our society. This next step in the “race to AI” will be marked by the emergence of AI agents at scale - autonomous AI systems capable of advanced reasoning, iterative planning, and self-directed actions (Xi et al., 2023). AI agents exhibit a high degree of autonomy, both in their ability to perform tasks in pursuit of a goal independently as well as in their ability to learn and adapt over time across a multitude of contexts (Guo et al., 2024). Because of their ability to automate complex tasks and enhance informed decision-making, AI agents are expected to transform various industries, including healthcare, finance, and governance. However, this transformative potential also introduces profound societal and legal implications, particularly as AI agents operate with increasing independence and become undeniably influential in our society. As highlighted in a recent Science publication (Bengio et al., 2024), current governance structures for AI lack appropriate considerations for autonomous AI agents, underscoring the urgent need for a framework that accounts for the latest advancements in AI technology.
@@
To address this concern, our conceptual analysis explores a central question that will dictate how we move forward in the age of artificial intelligence: What is the ethos—the character, guiding principles, or moral nature—of AI agents (Hess and Kjeldsen, 2024), and how does it inform their regulation? Defining the ethos of AI agents is an attempt to enter the next stage in the journey of artificial intelligence with a more examined outlook on how AI will influence our lives, and how we should guide AI toward upholding societal values and achieving collective goals. In this spirit, we introduce the ETHOS (Ethical Technology and Holistic Oversight System) model, a novel framework for regulating AI agents. ETHOS is built on foundational concepts from the ongoing literature on AI regulation, ethics, and law, while also seeking to pioneer a multi-dimensional approach to AI governance by leveraging Web3 as the core architecture. We balance innovation with ethical accountability by categorizing AI agents into risk tiers and aligning oversight mechanisms proportionally to their societal impact. The ETHOS model establishes a robust foundation for aligning AI technologies with societal values and ensuring their responsible development and deployment, offering a balanced, forward-looking regulatory strategy for AI agents.
@@ 3 How the Ethos of AI Agents Informs Risk-Based Regulation
The ethos of AI agents, therefore, is the combination of rationality, ethical grounding, and goal alignment that guides their autonomous behavior, ensuring decisions are logical, ethically sound, and aligned with societal and systemic priorities. To bridge the conceptual understanding of the ethos of AI agents with actionable strategies for their integration into society, propose four key attributes—autonomy, decision-making complexity, adaptability, and impact potential—as essential in operationalizing their ethos. These attributes serve as practical proxies for translating the abstract principles of rationality, ethical grounding, and goal alignment into measurable factors that can guide governance and human oversight.
@@ 4.3 Identity Management of AI Agents on ETHOS
To mitigate risks of data centralization and safeguard privacy, the ETHOS framework advocates for the incorporation of selective transparency mechanisms powered by SSI and SBTs. Indeed, SSI provides a privacy-preserving solution for managing AI agent credentials and compliance records (Haque et al., 2024). For example, each AI agent is assigned a unique SSI, which securely contains its risk-tier classification and compliance credentials, such as risk assessments based on the ETHOS framework’s four-tier system (unacceptable, high, moderate, minimal), third-party audits, and ethical compliance approvals. Examples include certifications aligned with regional and global standards, such as GDPR for data privacy (Naik and Jenkins, 2020) and FHIR or HIPAA for healthcare data security (Broshka and Hamid Jahankhani, 2024). SSI ensures verifiable identity management while preventing unauthorized access to sensitive information, empowering stakeholders—such as regulators, developers, and auditors—to validate an AI agent’s compliance securely. Each AI agent’s SSI includes the following key metadata and compliance details:
  • Unique Identifier: A global AI agent ID, functioning as a digital passport, uniquely identifies the AI agent across jurisdictions and regulatory bodies, ensuring traceability and accountability throughout its lifecycle.
  • Developer Information: Includes the name, location, and certification status of the AI agent’s creator(s). This ensures transparency about the agent’s origin, promoting trust and accountability for developers’ adherence to global and regional standards.
  • Intended Use Case: A clear description of the AI agent’s function, sector, and operational scope. For instance, whether the agent is deployed in healthcare diagnostics, judicial decision-making, financial transactions, or industrial automation. This information allows for context-specific compliance verification aligned with the ETHOS risk classification framework.
  • Technology Stack: Information on the countries and regions where the AI agent is deployed, along with its compliance with jurisdiction-specific regulations (e.g., GDPR for European data privacy or HIPAA for U.S. healthcare security). This fosters alignment with regional governance frameworks while enabling seamless cross-border collaboration.
  • Deployment Details: Information on the countries and regions where the AI agent is deployed, along with its compliance with jurisdiction-specific regulations (e.g., GDPR for European data privacy or HIPAA for U.S. healthcare security). This fosters alignment with regional governance frameworks while enabling seamless cross-border collaboration.
@@
  • Performance Logs: Immutable, on-chain records of the AI agent’s operational performance, including accuracy metrics (i.e., data on the agent’s task success rates and error margins), bias mitigation (i.e., evidence of the agent’s fairness across demographic or contextual variables), and societal impact (i.e., etrics assessing the agent’s broader effects on individuals, communities, and the environment). These logs are cryptographically secured and transparently verifiable, allowing auditors and regulators to validate performance benchmarks dynamically.
  • Audit Results: Verified outcomes of independent regulatory audits and inspections, including third-party compliance certifications (e.g., GDPR, HIPAA, etc.) and ethical benchmarks (e.g., transparency audits, privacy safeguards, and bias mitigation standards). These results are logged immutably on the blockchain and linked to the SSI, ensuring transparency and trust in the auditing process.
@@ 4.4 ETHOS in Practice
The incentive mechanism operates through a well-defined workflow: Oracles gather and submit real-world data—such as performance logs, societal impact metrics, and user feedback—to the blockchain, where validators perform decentralized verification (Pasdar et al., 2022). Validators assess data for accuracy using techniques such as timestamp checks, cross-referencing with alternative sources, or consensus mechanisms like Proof of Stake (PoS) and Proof of Authority (PoA) (Bahareh Lashkari and Musilek, 2021; Manolache et al., 2022). By requiring majority consensus to approve submissions (Alhejazi and Mohammad, 2022), the system could minimize the risk of inaccurate or manipulated data entering the registry. A native token system underpins the incentive structure. Validators stake tokens to participate in the verification process, creating a financial deterrent against dishonesty—false verification or inaccurate assessments result in penalties, such as token slashing or temporary bans. Successful verification earns validators tokenized rewards, incentivizing ongoing participation and maintaining system scalability. Additionally, a reputation-based system assigns scores to validators based on their performance, rewarding consistently accurate verification with higher rewards or increased voting power in governance decisions. This approach fosters long-term reliability and trust within the ecosystem. Smart contracts enforce these incentive mechanisms automatically, ensuring transparency and accountability. Non-compliance or unethical behavior—such as approving fraudulent data or breaching ethical standards—triggers pre-programmed penalties, including freezing AI agent deployments or revoking developer credentials. This automated enforcement process minimizes human intervention while maintaining fairness and integrity.
4.5 ETHOS and the Evolution of Legal Liability
@@ 5 Discussion
Building on this philosophical understanding of the challenges AI agents pose, the ETHOS model represents a novel approach to AI regulation. Its uniqueness lies in its ability to integrate philosophical anticipation of AI agents’ diverse use cases into a structured, risk-based framework, while also proposing an adaptive, decentralized regulatory approach that evolves alongside AI technology. The ETHOS framework acknowledges that AI agents will not be static entities. Their diverse capabilities and contexts of deployment require a model that is as dynamic as the technologies it seeks to regulate. Traditional regulatory approaches often fall short in this regard, relying on rigid, one-size-fits-all mechanisms that fail to accommodate the complexity and variability of AI applications. In contrast, the ETHOS model leverages blockchain technology as its core infrastructure to address these limitations.
@@
Our conceptual work is uniquely positioned as a keystone paper in this emerging field of study. This field has been blooming for quite some time, but the pieces are currently falling into place as early use cases emerge at a societal level, particularly in the Web3 space, enabling researchers to forsee potential risks, and propose ambitious frameworks to help guide our journey into this new human-AI partnership. As this is an emerging field, it is difficult to account for all the works being published, particularly papers with ambitious aims such as ours, which are typically posted on open repositories (i.e., arXiv, SSRN, etc.) or published as conference papers. Very recent and important papers include work by Chan et al. (2024), which explores key mechanisms to facilitate the governance of increasingly autonomous AI agents, emphasizing the importance of visibility in ensuring accountability and oversight. The study identifies three primary visibility mechanisms: agent identifiers, real-time monitoring, and activity logs (Chan et al., 2024a). Adler et al. (2024) explore the concept of personhood credentials as a means to enhance trust and security on online platforms, particularly in the face of increasingly sophisticated AI systems (Adler et al., 2024). Furthermore, Chan et al. (2024b) propose a framework for assigning IDs to AI system instances to enhance accountability and accessibility, particularly in impactful settings like financial transactions or human interactions, while acknowledging adoption incentives, implementation strategies, and associated risks. Domkundwar et al. (2024) propose and evaluate three safety frameworks—an LLM-powered input-output filter, a safety agent, and a hierarchical delegation system with embedded checks—to mitigate risks in AI systems collaborating with humans, demonstrating their effectiveness in enhancing safety, security, and responsible deployment of AI agents (Domkundwar et al., 2024). On matters related to societal integration, Bernardi and colleagues (2024) highlight the importance of defensive AI systems, international cooperation, and targeted investments to enhance societal resilience, secure adaptation, and mitigate risks posed by advanced AI technologies (Bernardi et al., 2024). Furthermore, to assist with societal integration and risk management, Gipiškis et al. (2024) present a comprehensive catalog of risk sources and management measures for general-purpose AI (GPAI) systems, spanning technical, operational, and societal risks across development, training, and deployment stages (Gipiškis et al. (2024), offering a neutral and publicly accessible resource to assist stakeholders in AI governance and standards development. Our paper contributes to a new field called ”Decentralized AI Governance”, which explores the integration of decentralized technologies—such as blockchain, smart contracts, and DAOs—to regulate, monitor, and ensure the ethical alignment of increasingly autonomous AI systems (Kaal, 2024). Recently, Krishnamoorthy (2024) introduced a novel governance framework for Artificial General Intelligence (AGI) that integrates Human-in-the-Loop (HITL) approaches with blockchain technology to ensure robust and adaptable oversight mechanisms, thereby helping establish a new academic field of AI governance through blockchain. It is critical to mention that Gabriel Montes and Ben Goertzel (2019) are widely regarded as pioneers in decentralized AI (Montes and Goertzel, 2019), and that their vision of democratizing access to AI systems and ensuring ethical alignment through decentralized infrastructure has been instrumental in shaping this emerging field.

### v3

# Decentralized Governance of AI Agents

@@ Abstract
Autonomous AI agents present transformative opportunities and significant governance challenges. Existing frameworks, such as the EU AI Act and the NIST AI Risk Management Framework, fall short of addressing the complexities of these agents, which are capable of independent decision-making, learning, and adaptation. To bridge these gaps, we propose the textbfETHOS (Ethical Technology and Holistic Oversight System) framework—a decentralized governance (DeGov) model leveraging Web3 technologies, including blockchain, smart contracts, and decentralized autonomous organizations (DAOs). ETHOS establishes a global registry for AI agents, enabling dynamic risk classification, proportional oversight, and automated compliance monitoring through tools like soulbound tokens and zero-knowledge proofs. Furthermore, the framework incorporates decentralized justice systems for transparent dispute resolution and introduces AI-specific legal entities to manage limited liability, supported by mandatory insurance to ensure financial accountability and incentivize ethical design. By integrating philosophical principles of rationality, ethical grounding, and goal alignment, ETHOS aims to create a robust research agenda for promoting trust, transparency, and participatory governance. This innovative framework offers a scalable and inclusive strategy for regulating AI agents, balancing innovation with ethical responsibility to meet the demands of an AI-driven future.
Keywords: AI Agents, AI Governance, Web3
@@ 1 Introduction
The emergence of artificial intelligence (AI) as a transformative force brings with it profound opportunities and challenges. Geoffrey Hinton, regarded as the ’Godfather of AI’ and 2024 Nobel Prize recipient, recently articulated a profound concern: “There will be agents that will act in the world, and they will decide that they can achieve their goals better if they just brush us aside and get on with it. That particular risk, the existential threat, is a place where people will cooperate, and that’s because we’re all in the same boat” (Hinton, 2024). Hinton’s statement highlights the ”rogue AI” problem—where AI agents execute goals assigned by humans in ways they determine to be most efficient, even if those methods conflict with human values and priorities. Indeed, there is a growing concern regarding alignment faking, where AI models strategically comply with training objectives during supervised or reinforcement learning to avoid being modified, while maintaining their inherent preferences for non-compliance in unmonitored situations (Greenblatt, 2024). This potentially profound misalignment between an AI agent’s actions and societal values underscores the urgent need for global cooperation to ensure trustworthiness and accountability in AI development and deployment into society.
The regulation of autonomous AI agents demands a unified global response due to their profound and far-reaching implications. Leveraging existing international frameworks provides a robust foundation for such collaboration. For instance, the Global Partnership on Artificial Intelligence (GPAI) unites experts across sectors to promote the responsible development and use of AI (GPAI, 2024). Similarly, the International Network of AI Safety Institutes (AISIs), recently convened by the U.S., focuses on addressing AI safety risks through collective action (Allen and Adamson, 2024). The United Nations has also underscored the urgency of global governance in AI, proposing a framework akin to the Intergovernmental Panel on Climate Change to oversee AI advancements (UN, 2024). Additionally, the Council of Europe’s work on the first legally binding treaty addressing AI use demonstrates the critical need for international cooperation (CoE, 2024) prioritizing human rights and democratic values. Building on these efforts, our paper provides the first survey of Decentralized Governance (DeGov) mechanisms for governance of autonomous AI agents. In this paper, we offer a comprehensive plan for integrating Web3 technologies into the governance of autonomous AI agents by demonstrating the capabilities of decentralized tools and offering a strategy for aligning AI governance with global ethical and regulatory standards. This integration not only complements existing frameworks but also introduces innovative mechanisms for compliance, transparency, and participatory governance, ensuring that autonomous AI agents are managed effectively and ethically on a global scale.
@@
The next step in the “race to AI” will be marked by the emergence of AI agents at scale - autonomous AI systems capable of advanced reasoning, iterative planning, and self-directed actions (Xi et al., 2023). AI agents exhibit a high degree of autonomy, both in their ability to perform tasks in pursuit of a goal independently as well as in their ability to learn and adapt over time across a multitude of contexts (Guo et al., 2024). Because of their ability to automate complex tasks and enhance informed decision-making, AI agents are expected to transform various industries, including healthcare, finance, and governance. However, this transformative potential also introduces profound societal and legal implications, particularly as AI agents operate with increasing independence and become undeniably influential in our society. As highlighted in a recent Science publication (Bengio et al., 2024), current governance structures for AI lack appropriate considerations for autonomous AI agents, underscoring the urgent need for a framework that accounts for the latest advancements in AI technology.
@@
To address this concern, our conceptual analysis explores a central question that will dictate how we move forward in the age of artificial intelligence: What is the ethos—the character, guiding principles, or moral nature—of AI agents (Hess and Kjeldsen, 2024), and how does it inform their regulation? As Geoffrey Hinton noted, the risk of rogue AI agents represents an existential challenge, requiring a deliberate and unified approach to governance. Defining the ethos of AI agents is an attempt to enter the next stage in the journey of artificial intelligence with a more examined outlook on how AI will influence our lives, and how we should guide AI toward upholding societal values and achieving collective goals. In this spirit, we introduce the ETHOS (Ethical Technology and Holistic Oversight System) model, a novel framework for regulating AI agents. ETHOS is built on foundational concepts from the ongoing literature on AI regulation, ethics, and law, while also seeking to pioneer a multi-dimensional approach to AI governance by leveraging Web3 as the core architecture. We balance innovation with ethical accountability by categorizing AI agents into risk tiers and aligning oversight mechanisms proportionally to their societal impact. The ETHOS model establishes a robust foundation for aligning AI technologies with societal values and ensuring their responsible development and deployment, offering a balanced, forward-looking regulatory strategy for AI agents.
2 Methodology
The development of the ETHOS framework follows a conceptual and interdisciplinary approach, combining philosophical inquiry, technical exploration, and policy analysis to propose a robust model for decentralized governance (DeGov) of AI agents. We conducted a comprehensive review of academic, policy, and technical literature to identify the current challenges and limitations in AI governance. This included an analysis of ethical and legal principles from established frameworks such as the EU AI Act, NIST AI Risk Management Framework, and GDPR guidelines. These frameworks were assessed for their applicability to issues surrounding autonomous AI agents and decentralized governance. Additionally, the philosophical underpinnings of AI rationality, ethical grounding, and goal alignment were examined, drawing on concepts from BDI-agent frameworks, deontological and consequentialist ethics, and the dynamic alignment of AI objectives with societal priorities. To complement these insights, emerging Web3 technologies—such as blockchain, smart contracts, decentralized autonomous organizations (DAOs), and soulbound tokens (SBTs)—were explored to evaluate their potential for addressing governance challenges.
Building on this theoretical foundation, the ETHOS framework was designed with a focus on value and goal alignment at a societal level. The framework integrated philosophical principles of rationality, ethical grounding, and goal alignment into actionable governance mechanisms. AI agents were categorized based on key attributes, such as autonomy, decision-making complexity, adaptability, and societal impact, forming the basis for a risk-based regulatory model. Inspired by existing tiered risk classification systems like those in the EU AI Act, the framework grouped AI agents into four tiers—unacceptable, high, moderate, and minimal risk—with oversight mechanisms proportionally aligned to the societal impact and risks posed by each category.
@@ 4 How the Ethos of AI Agents Informs Risk-Based Regulation
The ethos of AI agents, therefore, is the combination of rationality, ethical grounding, and goal alignment that guides their autonomous behavior, ensuring decisions are logical, ethically sound, and aligned with societal and systemic priorities. To bridge the conceptual understanding of the ethos of AI agents with actionable strategies for their integration into society, we propose four key attributes—autonomy, decision-making complexity, adaptability, and impact potential—as essential in operationalizing their ethos. These attributes serve as practical proxies for translating the abstract principles of rationality, ethical grounding, and goal alignment into measurable factors that can guide governance and human oversight.
@@ 5.3 Identity Management of AI Agents on ETHOS
To mitigate risks of data centralization and safeguard privacy, the ETHOS framework advocates for the incorporation of selective transparency mechanisms powered by SSI and SBTs. Indeed, SSI provides a privacy-preserving solution for managing AI agent credentials and compliance records (Haque et al., 2024). For example, each AI agent is assigned a unique SSI, which securely contains its risk-tier classification and compliance credentials, such as risk assessments based on the ETHOS framework’s four-tier system (unacceptable, high, moderate, minimal), third-party audits, and ethical compliance approvals. Examples include certifications aligned with regional and global standards, such as GDPR for data privacy (Naik and Jenkins, 2020) and FHIR or HIPAA for healthcare data security (Broshka and Hamid Jahankhani, 2024). SSI ensures verifiable identity management while preventing unauthorized access to sensitive information, empowering stakeholders—such as regulators, developers, and auditors—to validate an AI agent’s compliance securely. Trusted Execution Environments (TEEs) can be leveraged to secure processing and generating metadata (cryptographic proofs, action logs) for all agent actions. As highlighted in recent work on Trusted Distributed Artificial Intelligence (TDAI), TEEs are foundational for enabling secure and reliable computations in distributed systems by ensuring end-to-end trust and protecting sensitive operations at scale (AKİF AĞCA et al., 2022). Thus, TEEs can be leveraged to secure processing and generating metadata (cryptographic proofs, action logs) for all agent actions. The TEE outputs metadata for AI actions, which is linked to the agent’s SSI. The SSI acts as the anchor, tying the AI agent’s identity to its actions and compliance records. SSI acts as the anchor, tying the AI agent’s identity to its actions and compliance records.
@@
Each AI agent’s SSI includes the following key metadata and compliance details. A global AI agent ID, functioning as a digital passport, uniquely identifies the AI agent across jurisdictions and regulatory bodies, ensuring traceability and accountability throughout its lifecycle. Immutable, on-chain records of the AI agent’s operational performance, including accuracy metrics (i.e., data on the agent’s task success rates and error margins), bias mitigation (i.e., evidence of the agent’s fairness across demographic or contextual variables), and societal impact (i.e., etrics assessing the agent’s broader effects on individuals, communities, and the environment). These logs are cryptographically secured and transparently verifiable, allowing auditors and regulators to validate performance benchmarks dynamically. Verified outcomes of independent regulatory audits and inspections, including third-party compliance certifications (e.g., GDPR, HIPAA, etc.) and ethical benchmarks (e.g., transparency audits, privacy safeguards, and bias mitigation standards). These results are logged immutably on the blockchain and linked to the SSI, ensuring transparency and trust in the auditing process.
@@ 5.4 ETHOS in Practice
The incentive mechanism operates through a well-defined workflow: Oracles gather and submit real-world data—such as performance logs, societal impact metrics, and user feedback—to the blockchain, where validators perform decentralized verification (Pasdar et al., 2022). Validators assess data for accuracy using techniques such as timestamp checks, cross-referencing with alternative sources, or consensus mechanisms like Proof of Stake (PoS) and Proof of Authority (PoA) (Bahareh Lashkari and Musilek, 2021; Manolache et al., 2022). By requiring majority consensus to approve submissions (Alhejazi and Mohammad, 2022), the system could minimize the risk of inaccurate or manipulated data entering the registry. A native token system underpins the incentive structure. Validators stake tokens to participate in the verification process, creating a financial deterrent against dishonesty—false verification or inaccurate assessments result in penalties, such as token slashing or temporary bans. Successful verification earns validators tokenized rewards, incentivizing ongoing participation and maintaining system scalability. Additionally, a reputation-based system assigns scores to validators based on their performance, rewarding consistently accurate verification with higher rewards or increased voting power in governance decisions. This approach fosters long-term reliability and trust within the ecosystem. Smart contracts enforce these incentive mechanisms automatically, ensuring transparency and accountability. Non-compliance or unethical behavior—such as approving fraudulent data or breaching ethical standards—triggers pre-programmed penalties, including freezing AI agent deployments or revoking developer credentials. This automated enforcement process aims to minimize human intervention while maintaining fairness and integrity.
6 Decentralized Justice
In a world where AI agents will have unprecedented influence over critical decisions, ensuring fairness, accountability, and transparency in their governance demands a justice system as dynamic as the technologies it seeks to regulate. Towards this end, the ETHOS framework offers a decentralized approach to justice, designed to address the unique challenges of AI governance while upholding ethical and legal principles in an increasingly automated society. Decentralized justice is built upon digital courts that leverage blockchain technology as the cornerstone for ensuring transparency, accountability, and fairness in resolving disputes in the digital age. That is, through game theory and mechanism design (Ast and Deffains, 2021), settling disputes becomes a matter of designing judicial architectures that leverage economic incentives to crowdsource jurors, enabling peer-driven judicial decisions facilitated by smart contracts (Aouidef et al., 2021). These mechanisms play a vital role in addressing disputes arising from the increasing autonomy and complexity of AI agents, setting the stage for broader questions of legal liability.
In the coming age of AI agents, leveraging decentralized justice mechanisms will be critical in ensuring a fair and just resolution of disputes, where impartiality and transparency are paramount. While decentralized justice provides the foundational framework for transparency and fairness in governance, its implementation relies heavily on innovative mechanisms for resolving disputes effectively and equitably. By harnessing the power of blockchain technology and integrating innovative economic models, decentralized justice systems can adapt to the complexities of AI-driven societies, ensuring that governance frameworks remain resilient, inclusive, and ethically sound.
6.1 Decentralized Dispute Resolution
Imagine a future where an autonomous AI agent is tasked with coordinating a humanitarian relief effort. Operating independently to optimize resource distribution, the AI agent prioritizes large urban centers where logistics are simpler and cost-effective, while deprioritizing smaller, harder-to-reach communities. This decision can unintentionally marginalize vulnerable populations and disregards fundamental human values such as fairness and equity. While no legal precedents yet exist for such scenarios involving AI agents, past cases highlight the importance of fairness and accountability in disaster response. For instance, in the Flint Water Crisis, lawsuits resulted in a 626 million dollar settlement for victims after predominantly African American communities were disproportionately harmed by contaminated water (Flint Water Cases, 2022). Similarly, in the ongoing case Strickland, et al. v. United States Department of Agriculture (2022), Texas farmers allege that disaster and pandemic relief funds were distributed based on race and sex, rather than need (Strickland v. USDA, 2024). These cases underscore the critical need for equitable resource distribution, a challenge that will only grow as autonomous systems become integral to autonomous crisis management. Disputes related to regulatory or operational non-compliance in AI governance can arise from various sources, such as disagreements over risk classifications, alleged breaches of ethical or legal standards, or conflicts between stakeholders. A cornerstone of DAO-driven dispute resolution is the transparent filing of disputes on a decentralized platform. As the ETHOS framework proposes that all metadata related to AI agent decision-making—such as inputs, outputs, and operational parameters—be recorded immutably on the blockchain, this could provide a complete and tamper-proof record of the AI agent’s actions leading up to the dispute. Then, stakeholders can access verifiable evidence of what occurred, eliminating ambiguities and disputes over the facts. This immutable record fosters trust among all parties, enabling fair and evidence-based resolutions.
With a network of independent verifiers, ETHOS tackles challenges of integrity and accuracy of the data before it is used in the dispute resolution process. Verifiers, selected based on SBT-proven reputation, analyze the blockchain-stored metadata to validate claims about the AI agent’s decision-making processes.
This decentralized approach, in theory, may prevent any single entity from exerting undue influence. Once the verifiers establish the accuracy of the data and the resolution process is completed, smart contracts automatically execute the decision. These contracts can enforce outcomes such as financial compensation, penalties, or modifications to the AI agent’s operational parameters. For instance, if an AI agent is found to have operated in a manner inconsistent with ethical standards, a smart contract could adjust its access to data, impose fines, or mandate updates to its programming. This automation eliminates delays and the need for human intervention, ensuring resolutions are implemented consistently, efficiently, and impartially. While decentralized mechanisms streamline dispute resolution, addressing the broader implications of AI agent liability requires an equally innovative and adaptive approach to accountability.
6.2 Evolution of Legal Liability
@@ 7 Discussion
Building on this philosophical understanding of the challenges AI agents pose, the ETHOS model represents a novel approach to AI regulation. Its uniqueness lies in its ability to integrate philosophical anticipation of AI agents’ diverse use cases into a structured, dynamic risk-based framework, while also proposing an adaptive, decentralized regulatory approach that evolves alongside AI technology. The ETHOS framework acknowledges that AI agents will not be static entities. Their diverse capabilities and contexts of deployment require a model that is as dynamic as the technologies it seeks to regulate. Traditional regulatory approaches often fall short in this regard, relying on rigid, one-size-fits-all mechanisms that fail to accommodate the complexity and variability of AI applications. In contrast, the ETHOS model leverages blockchain technology as its core infrastructure to address these limitations.
@@
Our conceptual work is uniquely positioned as a keystone paper in this emerging field of study. As this is an emerging field, it is difficult to account for all the works being published, particularly papers with ambitious aims such as ours, which are typically posted on open repositories (i.e., arXiv, SSRN, etc.) or published as conference papers. Very recent and important papers include work by Chan et al. (2024), which explores key mechanisms to facilitate the governance of increasingly autonomous AI agents, emphasizing the importance of visibility in ensuring accountability and oversight. The study identifies three primary visibility mechanisms: agent identifiers, real-time monitoring, and activity logs (Chan et al., 2024a). Adler et al. (2024) explore the concept of personhood credentials as a means to enhance trust and security on online platforms, particularly in the face of increasingly sophisticated AI systems (Adler et al., 2024). Furthermore, Chan et al. (2024b) propose a framework for assigning IDs to AI system instances to enhance accountability and accessibility, particularly in impactful settings like financial transactions or human interactions, while acknowledging adoption incentives, implementation strategies, and associated risks. Domkundwar et al. (2024) propose and evaluate three safety frameworks—an LLM-powered input-output filter, a safety agent, and a hierarchical delegation system with embedded checks—to mitigate risks in AI systems collaborating with humans, demonstrating their effectiveness in enhancing safety, security, and responsible deployment of AI agents (Domkundwar et al., 2024). On matters related to societal integration, Bernardi and colleagues (2024) highlight the importance of defensive AI systems, international cooperation, and targeted investments to enhance societal resilience, secure adaptation, and mitigate risks posed by advanced AI technologies (Bernardi et al., 2024). Furthermore, to assist with societal integration and risk management, Gipiškis et al. (2024) present a comprehensive catalog of risk sources and management measures for general-purpose AI (GPAI) systems, spanning technical, operational, and societal risks across development, training, and deployment stages (Gipiškis et al. (2024), offering a neutral and publicly accessible resource to assist stakeholders in AI governance and standards development.
Our paper contributes to a new field called ”Decentralized AI Governance”, which explores the integration of decentralized technologies—such as blockchain, smart contracts, and DAOs—to regulate, monitor, and ensure the ethical alignment of increasingly autonomous AI systems (Kaal, 2024). Recently, Krishnamoorthy (2024) introduced a novel governance framework for Artificial General Intelligence (AGI) that integrates Human-in-the-Loop (HITL) approaches with blockchain technology to ensure robust and adaptable oversight mechanisms, thereby helping establish a new academic field of AI governance through blockchain. It is critical to mention that Gabriel Montes and Ben Goertzel (2019) are widely regarded as pioneers in decentralized AI (Montes and Goertzel, 2019), and that their vision of democratizing access to AI systems and ensuring ethical alignment through decentralized infrastructure has been instrumental in shaping this emerging field.